OpenStudy (anonymous):

√a³b^5c² simplify and show work

OpenStudy (anonymous):

\[\sqrt{a ^{3}b ^{5} c ^{2}}\] =\[\sqrt{a ^{2}ab ^{4}bc ^{2}}\] =\[ab ^{2}c \sqrt{ab}\]